# AI in Supply Chain Risk Assessment: A Systematic Literature Review and Bibliometric Analysis

### v4

@@ Abstract
Supply chain risk assessment (SCRA) has witnessed a profound evolution through the integration of artificial intelligence (AI) and machine learning (ML) techniques, revolutionizing predictive capabilities and risk mitigation strategies. The significance of this evolution stems from the critical role of robust risk management strategies in ensuring operational resilience and continuity within modern supply chains. Previous reviews have outlined established methodologies but have overlooked emerging AI/ML techniques, leaving a notable research gap in understanding their practical implications within SCRA. This paper conducts a systematic literature review combined with a comprehensive bibliometric analysis. We meticulously examined 1,439 papers and derived key insights from 51 articles published between 2015 and 2024. The review fills this research gap by addressing pivotal research questions and exploring existing AI/ML techniques, methodologies, findings, and future trajectories, thereby providing a more encompassing view of the evolving landscape of SCRA. Our study unveils the transformative impact of AI/ML models, such as Random Forest, XGBoost, and hybrids, in substantially enhancing precision within SCRA. It underscores adaptable post-COVID strategies, advocating for resilient contingency plans and aligning with evolving risk landscapes. Significantly, this review surpasses previous examinations by accentuating emerging AI/ML techniques and their practical implications within SCRA. Furthermore, it highlights the contributions through a comprehensive bibliometric analysis, revealing publication trends, influential authors, and highly cited articles. This analysis provides an encompassing view of the field’s growth, offering insights into its evolution. This comprehensive review provides a roadmap for practitioners and researchers, offering insights into fortifying SCRM strategies through AI integration, ultimately contributing to a deeper understanding of evolving trends and applications in this dynamic field.
@@ 1 Introduction
Supply chain management (SCM) has long grappled with various challenges from various sources. Past outbreaks of infectious diseases, geological upheavals like earthquakes, and other natural catastrophes have put supply chains (SCs) at risk, albeit on a limited scale (Govindan et al.,, 2020). These incidents have illustrated how interrelationships within SCs have formed intricate risk contagion networks, making them susceptible to contagion effects (Agca et al.,, 2022). The cascade effect, characterized by risk spillover from one enterprise to another, exacerbates the challenges, culminating in a chain reaction of SC Resilience (SCR) disasters (Roukny et al.,, 2018).
Across the globe, companies confront formidable challenges across all stages of their SCs. Suppliers failing to meet delivery obligations, unpredictable shifts in customer demands, and episodes of panic buying are just a few examples of the hurdles companies face (Ivanov,, 2020). In pursuing market leadership, enterprises must also contend with the complexity introduced by the management and launch of innovation projects, further complicating SCM (Kwak et al.,, 2018). Furthermore, digital transformation has ushered in a wave of technological advancements in SC operations (Kwak et al.,, 2018).
@@
However, the most significant disruption in recent times emerged during the COVID-19 pandemic, impacting global SCs profoundly. The disruptions encompassed not only the movement of people but also the flow of raw materials and finished goods, along with extensive disruptions in factory and SC operations (Sheng et al.,, 2021). These disruptions ushered in unprecedented challenges for SC professionals as they grappled with an entirely new reality (Araz et al.,, 2020; Craighead et al.,, 2020). The pandemic represents a distinctive form of SC disruption, distinct from natural or man-made disasters (e.g., the Japanese earthquake and 9/11 attacks) and disruptions driven by evolving technologies and changing customer attitudes (Zhang et al.,, 2020).
While existing guidance exists for predicting, managing, and responding to these disruptions, the challenges brought about by COVID-19 have underscored the paramount importance of risk management, unlike any previous disruption (Bode et al.,, 2011; Craighead et al.,, 2020). Consequently, the concept of SCR in the context of COVID-19 has emerged as a critical area necessitating further exploration and development.
@@
SCs are not the sole victims in this precarious environment, as the ripple effect extends to upstream and downstream enterprises. Practical risk management programs are imperative for enterprises to avert SCR and enhance SC risk management (SCRM). Moreover, incorporating advanced AI technologies, such as ML, can provide predictive capabilities for SCRM. ML algorithms have showcased their ability to identify abnormal risk factors and derive predictive insights from historical data (Guo et al.,, 2021; Mohanty et al.,, 2021). By harnessing ML, enterprises can detect risk factors and anticipate market demands and potential risk scenarios (Punia et al.,, 2020; Wu et al., 2022a, ). ML’s proficiency in handling non-linear relationships further bolsters its superiority over traditional linear models. Additionally, its aptitude for processing unstructured data, a challenge where traditional models falter, positions ML as a formidable tool for addressing time, cost, and resource constraints within SCs.
However, despite its immense potential, SC researchers have historically exhibited limited familiarity with ML compared to other SCM aspects like mathematical programming and stochastic optimization (Liu et al.,, 2019; Shahed et al.,, 2021). Furthermore, the classification of ML algorithms remains unclear (Janiesch et al.,, 2021; Xu and Jackson,, 2019). Bridging this knowledge gap and exploring the value of ML for SCRM through interdisciplinary integration is a critical research need. Remarkably, there has been no prior effort to scrutinize the SCRM literature within an ML environment, in stark contrast to previous reviews primarily centered on risk definition, classification, and management strategies.
@@
Companies should augment their analytical capabilities by harnessing organizational knowledge to enhance SCR, thereby elevating their information capabilities (Wong et al.,, 2020). As underscored by previous studies, the role of AI-enabled technologies extends to promoting innovations for enhanced SC performance (SCP) (Baryannis et al., 2019b, ; Nayal et al.,, 2021). The adaptation capabilities and information processing prowess offered by AI techniques hold the potential to enhance SCP (Belhadi et al., 2021a, ). Notably, AI has found application across diverse sectors, improving flexibility and communication and reducing undue fluctuations for successful project execution (Lalmi et al.,, 2021). To navigate the impact of risks and disruptions, (Katsaliaki et al.,, 2022) advocates integrating three critical facets: long-term partnerships, IT applications for business enhancement, and government policies that facilitate adaptability.
@@
The rest of the article is organized as follows: the “Literature Review” examines the current body of knowledge. The “Methodology” section details the study’s design, including search strategies and selection criteria. The “Bibliometric Analysis” provides insights from bibliometric tools, with visualizations and observations on AI in SCRA. A structured framework for SCRA is given in “Framework for SCRA using AI section. In the “Content Analysis of AI Techniques in SCRA,” the review discusses various AI techniques, highlighting their benefits, limitations, and real-world applications. Practical insights for industry are provided in the “Practical and Managerial Implications” section, offering actionable guidance for practitioners. The review also outlines key “Challenges and Limitations” in this area, and suggests “Future Research Directions” for ongoing exploration. The paper concludes with a summary of findings and recommendations in the “Conclusion.”
@@ 2 Literature Review
Deiva Ganesh and Kalpana, (2022) conducted a comprehensive and descriptive literature study to identify AI and ML approaches in SCRM stages. This study examined SCRM research publications from three scientific databases from 2010 to 2021. They proposed a data analytics, simulation, and optimization framework that might produce a comprehensive SC risk identification, assessment, mitigation, and monitoring strategy. Hence, leveraging AI, Blockchain, and the Industrial Internet of Things (IIoT) to create smarter SCs can transform how firms handle uncertainty. This analysis of Deiva Ganesh and Kalpana, (2022) does not evaluate blockchain, big data, and IIoT-related SC articles, which may limit information. Due to limited research, only recent English-language publications on SCRM and AI were included. Nimmy et al., (2022) did a thorough literature evaluation on operational risk assessment methods and whether AI explains them. The report suggested that risk managers use auditable SC operational risk management methodologies to understand why they should take a risk management action rather than just what to do. While they employed the SLR method, they could only examine some AI solutions for SCRM. They suggested hybridizing SC operational risk management and explainable AI (XAI) to obtain XAI-like characteristics in future research. Li et al., (2023) analyzed COVID-19 SCR research history, present, and future. In particular, supervised ML classifies 1717 SCR papers into 11 subject categories. Each cluster was then studied in the context of COVID-19, indicating three related skills (interconnectedness, transformability, and sharing) on which enterprises could work to develop a more resilient SC post-COVID. Their data only came from Scopus’ core collection, which might affect the results; thus, they suggested adding WoS and EBSCO to the evaluation (Xu et al.,, 2020). Given the fast expansion of SCR research, they only picked English-language articles, which may have excluded valuable knowledge. They recommended network analysis to determine cluster linkages and SCR literary themes.
@@ 3.1 Study Design
The SLR followed a robust methodology, adhering to the guidelines proposed by Denyer and Tranfield, (2009). This approach aimed to identify and analyze research articles focusing on integrating AI/ML techniques within SCRA.
@@ 3.2 Search Strategy
We conducted a comprehensive search using two major academic databases: Google Scholar and Web of Science (WoS). We used the “Publish or Perish 8” (Harzing,, 2016) software for the Google Scholar search to filter articles published between 2015 and 2024. The initial search yielded 1000 articles. In the WoS search, we employed the following search strings to retrieve relevant articles:
@@
Date range 2015–2024 Focuses on recent studies, ensuring relevance to current advancements in AI for SCRA.
@@ 3.3 Article Selection Process
To ensure the selection of high-quality and relevant studies, we followed a systematic, multi-stage screening process guided by the PRISMA (Preferred Reporting Items for Systematic Reviews and Meta-Analyses) framework. The selection process was designed to minimize potential bias and ensure transparency at each stage. The stages are detailed as follows:
@@
    Removal of Duplicates: To avoid redundant analysis, we first identified and removed duplicate articles between the two databases, excluding 360 duplicate entries. This step is standard in systematic reviews to ensure each study contributes uniquely to the analysis.
@@
    Filtering by Publisher Reliability: We excluded 398 articles published by sources identified as less reliable, focusing only on those published in reputable, peer-reviewed journals. This criterion helps maintain high academic integrity and ensures that the final selection is based on studies that meet established research quality standards.
@@
  5. 5.
    Checking for Relevance (Title, Abstract, and Conclusion Screening): The remaining articles were reviewed based on their title, abstract, and conclusions to assess their relevance to the research questions of this study. Articles that did not focus specifically on AI applications in SCRA were excluded. This step eliminated an additional 383 articles, leaving a final sample of 51 articles for in-depth review.
@@ 3.5 Final Selection
After applying all criteria, 51 articles were deemed relevant and of high quality for inclusion in this review. The selected articles cover various AI methodologies applied to different SC risk scenarios, providing a comprehensive foundation for our analysis. According to Figure 1, 434 articles were included in the bibliometric analysis to provide a broad understanding of the field, considering various quantitative metrics. The 51 shortlisted articles represent a more focused subset. We selected them based on additional qualitative criteria, such as relevance, methodology, and contribution to the research.
@@ 3.6 Data Extraction and Synthesis
The selected articles underwent detailed analysis, focusing on methodologies, AI/ML techniques employed, key findings, and implications for SCRA. Data extraction forms were filled in to gather pertinent information for synthesis and further analysis. To understand the research landscape comprehensively, we performed bibliometric and cocitation analysis of the 434 articles resulting from the filtering process. The bibliometric analysis helped us identify the field’s most influential authors, journals, and key research themes. Cocitation analysis allowed us to identify clusters of related articles and determine their interconnections. Our study has several limitations that should be acknowledged. First, the search was limited to articles published between 2015 and 2024, and it is possible that relevant articles published before or after this timeframe were not included.
@@ 4.1 Publication Trend Over Time
Figure 2: Annual publication trend (2015-2024) of the reviewed literature.
Figure 2 illustrates the annual publication trends from 2015 to 2024. A significant rise in publications is observed post-2020, likely influenced by the COVID-19 pandemic, highlighting the critical need for resilient SC systems. The peak in research output during 2021 and 2022 reflects the heightened importance of AI/ML-driven risk assessment in tackling complex SC challenges. However, the slight decline in subsequent years may indicate a maturing research field, transitioning from exploratory studies to more application-oriented implementations in industry.

### v5

@@ Abstract
Supply chain risk assessment (SCRA) is pivotal for ensuring resilience in increasingly complex global supply networks. While existing reviews have explored traditional methodologies, they often neglect emerging artificial intelligence (AI) and machine learning (ML) applications and mostly lack combined systematic and bibliometric analyses. This study addresses these gaps by integrating a systematic literature review with bibliometric analysis, examining 1,903 articles (2015–2025) from Google Scholar and Web of Science, with 54 studies selected through PRISMA guidelines. Our findings reveal that ML models, including Random Forest, XGBoost, and hybrid approaches, significantly enhance risk prediction accuracy and adaptability in post-pandemic contexts. The bibliometric analysis identifies key trends, influential authors, and institutional contributions, highlighting China and the United States as leading research hubs. Practical insights emphasize the integration of explainable AI (XAI) for transparent decision-making, real-time data utilization, and blockchain for traceability. The study underscores the necessity of dynamic strategies, interdisciplinary collaboration, and continuous model evaluation to address challenges such as data quality and interpretability. By synthesizing AI-driven methodologies with resilience frameworks, this review provides actionable guidance for optimizing supply chain risk management, fostering adaptability, and informing future research in evolving risk landscapes.
@@ 1 Introduction
Supply chain management (SCM) has long grappled with various challenges from various sources. Past outbreaks of infectious diseases, geological upheavals like earthquakes, and other natural catastrophes have put supply chains (SCs) at risk, albeit on a limited scale (Govindan et al.,, 2020). These incidents have illustrated how interrelationships within SCs have formed intricate risk contagion networks, making them susceptible to contagion effects (Agca et al.,, 2022). The cascade effect, characterized by risk spillover from one enterprise to another, exacerbates the challenges, culminating in a chain reaction of SC Resilience (SCR) disasters (Roukny et al.,, 2018). Across the globe, companies confront formidable challenges across all stages of their SCs. Suppliers failing to meet delivery obligations, unpredictable shifts in customer demands, and episodes of panic buying are just a few examples of the hurdles companies face (Ivanov,, 2020). In pursuing market leadership, enterprises must also contend with the complexity introduced by the management and launch of innovation projects, further complicating SCM (Kwak et al.,, 2018). Furthermore, digital transformation has ushered in a wave of technological advancements in SC operations (Kwak et al.,, 2018).
@@
However, the most significant disruption in recent times emerged during the COVID-19 pandemic, impacting global SCs profoundly. The disruptions encompassed not only the movement of people but also the flow of raw materials and finished goods, along with extensive disruptions in factory and SC operations (Sheng et al.,, 2021). These disruptions ushered in unprecedented challenges for SC professionals as they grappled with an entirely new reality (Araz et al.,, 2020; Craighead et al.,, 2020). The pandemic represents a distinctive form of SC disruption, distinct from natural or man-made disasters (e.g., the 2011 earthquake in Japan and the 9/11 attacks) and disruptions driven by evolving technologies and changing customer attitudes (Zhang et al.,, 2020). While existing guidance exists for predicting, managing, and responding to these disruptions, the challenges brought about by COVID-19 have underscored the paramount importance of risk management, unlike any previous disruption (Bode et al.,, 2011; Craighead et al.,, 2020). Consequently, the concept of SCR in COVID-19 has emerged as a critical area necessitating further exploration and development.
@@
SCs are not the sole victims in this precarious environment, as the ripple effect extends to upstream and downstream enterprises. Practical risk management programs are imperative for enterprises to avert SCR and enhance SC risk management (SCRM). AI encompasses various techniques that enable autonomous decision-making in complex and uncertain environments. As noted by Baryannis et al., 2019b , an SCRM approach is considered artificially intelligent if it can autonomously decide on a course of action that leads to success in SCRM-related objectives while operating under a partially unknown SC environment. The AI techniques utilized in this context span from traditional symbolic AI, which relies on mathematical or knowledge-based problem representations, to sub-symbolic AI, including fuzzy systems and evolutionary computation, and statistical AI, which encompasses ML approaches. Among these, ML algorithms have showcased their ability to identify abnormal risk factors and derive predictive insights from historical data (Guo et al.,, 2021; Mohanty et al.,, 2021). By harnessing ML, enterprises can detect risk factors and anticipate market demands and potential risk scenarios (Punia et al.,, 2020; Wu et al., 2022a, ). ML’s proficiency in handling non-linear relationships further bolsters its superiority over traditional linear models. Additionally, its aptitude for processing unstructured data, a challenge where traditional models falter, positions ML as a formidable tool for addressing time, cost, and resource constraints within SCs. ML has emerged as the most suitable technique for risk assessment in SC. Therefore, this review focuses specifically on ML-based approaches in the context of SC resilience assessment.
@@
However, despite its immense potential, SC researchers have historically exhibited limited familiarity with ML compared to other SCM aspects like mathematical programming and stochastic optimization (Liu et al.,, 2019; Shahed et al.,, 2021). Furthermore, the classification of ML algorithms remains unclear (Janiesch et al.,, 2021; Xu and Jackson,, 2019). Bridging this knowledge gap and exploring the value of ML for SCRM through interdisciplinary integration is a critical research need. Remarkably, there has been no prior effort to scrutinize the SCRM literature within an ML environment, in stark contrast to previous reviews centered on risk definition, classification, and management strategies. Companies should augment their analytical capabilities by harnessing organizational knowledge to enhance SCR, thereby elevating their information capabilities (Wong et al.,, 2020). As underscored by previous studies, the role of AI-enabled technologies extends to promoting innovations for enhanced SC performance (SCP) (Baryannis et al., 2019b, ; Nayal et al.,, 2021). The adaptation capabilities and information processing prowess offered by AI techniques hold the potential to enhance SCP (Belhadi et al., 2021a, ). Notably, AI has found application across diverse sectors, improving flexibility and communication and reducing undue fluctuations for successful project execution (Lalmi et al.,, 2021). To navigate the impact of risks and disruptions, (Katsaliaki et al.,, 2022) advocates integrating three critical facets: long-term partnerships, IT applications for business enhancement, and government policies that facilitate adaptability.
@@
The rest of the article is organized as follows: the “Literature Review” examines the current body of knowledge. The “Methodology” section details the study’s design, including search strategies and selection criteria. The “Bibliometric Analysis” provides insights from bibliometric tools, with visualizations and observations on AI in SCRA. A structured framework for SCRA is given in “Framework for SCRA using AI section. In the “Content Analysis of AI Techniques in SCRA,” the review discusses various AI techniques, highlighting their benefits, limitations, and real-world applications. Practical insights for the industry are provided in the “Practical and Managerial Implications” section, offering actionable guidance for practitioners. The review also outlines key “Challenges and Limitations” in this area, and suggests “Future Research Directions” for ongoing exploration. The paper concludes with a summary of findings and recommendations in the “Conclusion.”
@@ 2 Literature Review
Deiva Ganesh and Kalpana, (2022) conducted a comprehensive and descriptive literature study to identify AI and ML approaches in SCRM stages. This study examined SCRM research publications from three scientific databases from 2010 to 2021. They proposed a data analytics, simulation, and optimization framework that might produce a comprehensive SC risk identification, assessment, mitigation, and monitoring strategy. Hence, leveraging AI, Blockchain, and the Industrial Internet of Things (IIoT) to create smarter SCs can transform how firms handle uncertainty. This analysis of Deiva Ganesh and Kalpana, (2022) does not evaluate blockchain, big data, and IIoT-related SC articles, which may limit information. Due to limited research, only recent English-language publications on SCRM and AI were included. Nimmy et al., (2022) did a thorough literature evaluation on operational risk assessment methods and whether AI explains them. The report suggested that risk managers use auditable SC operational risk management methodologies to understand why they should take a risk management action rather than just what to do. While they employed the SLR method, they could only examine some AI solutions for SCRM. They suggested hybridizing SC operational risk management and explainable AI (XAI) to obtain XAI-like characteristics in future research. Li et al., (2023) analyzed COVID-19 SCR research history, present, and future. In particular, supervised ML classifies 1717 SCR papers into 11 subject categories. Each cluster was then studied in the context of COVID-19, indicating three related skills (interconnectedness, transformability, and sharing) on which enterprises could work to develop a more resilient SC post-COVID. Their data only came from Scopus’ Core Collection, which might affect the results; thus, they suggested adding WoS and EBSCOhost to the evaluation. (Xu et al.,, 2020). Given the fast expansion of SCR research, they only picked English-language articles, which may have excluded valuable knowledge. They recommended network analysis to determine cluster linkages and SCR literary themes.
@@ 3.1 Study Design
Our approach is guided by the structured methodology of Denyer and Tranfield, (2009) for systematic literature reviews, ensuring rigor in study identification and synthesis. Additionally, we adhered to the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) framework to enhance transparency and minimize bias in the study selection process. This approach aimed to identify and analyze research articles on integrating AI/ML techniques within SCRA.
@@ 3.2 Search Strategy
We conducted a comprehensive search using two major academic databases: Google Scholar and Web of Science (WoS) Core Collection. The WoS Core Collection, recognized as a high-quality digital literature database, has gained widespread acceptance among researchers and is currently regarded as one of the most appropriate databases for bibliometric analysis (Wang et al.,, 2024). Compared to other databases, WoS offers more extensive data coverage and provides data formats compatible with analysis tools such as VOSviewer. We used the “Publish or Perish 8” (Harzing,, 2016) software for the Google Scholar search to extract all data of the most relevant articles suitable for bibliometric analysis. The initial search on Google Scholar using Publish or Perish yielded 1000 articles, and WoS, using advanced search, yielded 903 articles before specifying any date range. We employed the following search strings to retrieve relevant articles:
@@
Date range 2015–2025 Focuses on recent studies, ensuring relevance to current advancements in AI for SCRA.
@@
Unrelated methodologies Exclusion Removes studies that employ computational models or frameworks that do not align with the research scope.
Limited generalizability Exclusion Excludes studies focusing on highly specific contexts that cannot be applied to broader AI-driven SCRA applications.
@@ 3.3 Article Selection Process
To ensure the selection of high-quality and relevant studies, we followed a systematic, multi-stage screening process guided by the PRISMA framework. The selection process was designed to minimize potential bias and ensure transparency at each stage. The stages are detailed as follows:
@@
    Date Range Filtering: We restricted our review to studies published between 2015–2025, excluding 23 articles, ensuring the inclusion of recent and relevant research. This step ensures that each study contributes uniquely to the analysis and aligns with the latest advancements in the field.
@@
    Removal of Duplicates: To avoid redundant analysis, we first identified and removed duplicate articles between the two databases, excluding 451 duplicate entries. This step is standard in systematic reviews to ensure each study contributes uniquely to the analysis.
@@
    Filtering by Publisher Reliability: We excluded 398 articles published by sources identified as less reliable, focusing only on those published in reputable publishers (IEEE, Springer Nature, Elsevier, Taylor & Francis, MDPI, Sage, Wiley, Nature Portfolio, ACM, etc.). This criterion helps maintain high academic integrity and ensures the final selection is based on studies that meet established research quality standards.
  4. 4.
@@
  6. 6.
    Checking for Relevance (Title, Abstract, and Conclusion Screening): The remaining articles were reviewed based on their title, abstract, and conclusions to assess their relevance to the research questions of this study. Articles that did not focus specifically on AI applications in SCRA were excluded. This step eliminated 515 articles.
  7. 7.
    Exclusion of Studies with Unrelated Methodologies: After relevance screening, we further analyzed the methodologies used in the remaining studies. Articles that employed computational models or frameworks that did not align with our scope were removed. This step excluded 128 articles, ensuring the final selection focused on studies directly relevant to the research objectives.
  8. 8.
    Exclusion of Studies with Limited Generalizability: We excluded 81 highly context-specific studies that lacked generalizability to broader AI-driven SCRA applications, leaving a final sample of 54 articles for in-depth review. These studies primarily focused on niche use cases that do not translate well to wider SCRA applications.
@@ 3.5 Final Selection
After applying all criteria, 54 articles were deemed relevant and of high quality for inclusion in this review. The selected articles cover various AI methodologies applied to different SC risk scenarios, providing a comprehensive foundation for our analysis. According to Figure 1, 778 articles were included in the bibliometric analysis to provide a broad understanding of the field, considering various quantitative metrics. The 54 shortlisted articles represent a more focused subset. We selected them based on additional qualitative criteria, such as relevance, methodology, and contribution to the research.
@@ 3.6 Data Extraction and Synthesis
The selected articles underwent detailed analysis, focusing on methodologies, AI/ML techniques employed, key findings, and implications for SCRA. Data extraction forms were filled in to gather pertinent information for synthesis and further analysis. To understand the research landscape comprehensively, we performed bibliometric and cocitation analysis of the 778 articles resulting from the filtering process. The bibliometric analysis helped us identify the field’s most influential authors, journals, and key research themes. Cocitation analysis allowed us to identify clusters of related articles and determine their interconnections. Our study has several limitations that should be acknowledged. First, the search was limited to articles published between 2015 and 2025, and it is possible that relevant articles published before or after this timeframe were not included.
@@ 4.1 Publication Trend Over Time
Figure 2: Annual publication trend (2015-2025) of the 778 published literature on AI in SCRA topic.
Figure 2 illustrates the annual publication trends from 2015 to 2025. A significant rise in publications is observed post-2020, likely influenced by the COVID-19 pandemic, highlighting the critical need for resilient SC systems. The peak in research output during 2024 reflects the heightened importance of AI/ML-driven risk assessment in tackling complex SC challenges. However, the slight decline in 2025 is because of the first quarter of the year, which is expected to peak following the exponential trend line.